OpenStudy (kl0723):

Limit of the following function as n approaches zero:

OpenStudy (kl0723):

\[nsin(\frac{ 1 }{ n })=?\]

OpenStudy (kl0723):

n approaches infinity''' can somebody help me understand this problem?

myininaya (myininaya):

big hint :) \[-1 \le \sin(\frac{1}{n} )\le 1\]

myininaya (myininaya):

oh are you doing two different problems?

myininaya (myininaya):

I see n->0 in the title and n->infty in your last post

OpenStudy (kl0723):

I meant to say infinity

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}\frac{\sin(\frac{1}{n})}{\frac{1}{n}}\] if n->infty then 1/n->?

myininaya (myininaya):

\[\lim_{u \rightarrow 0}\frac{\sin(u)}{u}=?\]

OpenStudy (kl0723):

zero, the fraction becomes infinitely small

myininaya (myininaya):

so as u->0 then sin(u)/u->?

myininaya (myininaya):

there is a famous limit proved in many calculus books

OpenStudy (kl0723):

sin(u)/(u) would yield 0 over 0

myininaya (myininaya):

well i mean sounds like you want to do l'hospital now instead of just remembering by sandwish theorem we have sin(u)/u->1 as u->0

myininaya (myininaya):

 so we have the following inequality \[\text{ area of the smaller triangle } \le \text{ area of sector } \le \text{ area of bigger triangle }\] \[\frac{1}{2} \sin(\theta) \cos(\theta) \le \frac{ \theta}{2} \le \frac{1}{2} \tan(\theta) \\ \sin(\theta) \cos(\theta) \le \theta \le \frac{\sin(\theta)}{\cos(\theta)} \\ \text{ assume } \sin(\theta)>0 \\ \text{ dividing both sides by } \sin(\theta) \\ \cos(\theta) \le \frac{ \theta}{ \sin(\theta)} \le \frac{1}{\cos(\theta)} \\ \lim_{\theta \rightarrow 0^+}\frac{ \sin(\theta)}{\theta}=1\] we can show in a similar way from the left we also have 1